OpenStudy (theedwardsfamily):

Variable y varies directly with variable x, and y = 18 when x = 4. What is the constant of variation in decimal form?

OpenStudy (theedwardsfamily):

@confluxepic

OpenStudy (theedwardsfamily):

help...

OpenStudy (theedwardsfamily):

please...

OpenStudy (theedwardsfamily):

@Destinymasha @adajiamcneal @Holly00d1248 @Legends @Luigi0210 @paki @oOKawaiiOo @snazzychazzy @TheSmartOne

OpenStudy (anonymous):

I AM HERE

OpenStudy (anonymous):

ok i know what to do

OpenStudy (anonymous):

so to do this you \[y=xk\]

OpenStudy (anonymous):

so its

OpenStudy (theedwardsfamily):

ok

OpenStudy (anonymous):

18=4k

OpenStudy (anonymous):

divide bothe sides by 4

OpenStudy (anonymous):

so the answer is 4.5

OpenStudy (anonymous):

i hope i helped :)

OpenStudy (theedwardsfamily):

yea thats wat i got

OpenStudy (theedwardsfamily):

you did thnx

OpenStudy (anonymous):

if i did plz medal

OpenStudy (anonymous):

np

OpenStudy (ybarrap):

$$ y={\alpha x}\\ \alpha ={y\over x}\\ \alpha ={18\over 4}\\ $$ Where \(\alpha \) is your constant of proportionality